### v1

# FANformer: Improving Large Language Models Through Effective Periodicity Modeling

Abstract
Periodicity, as one of the most important basic characteristics, lays the foundation for facilitating structured knowledge acquisition and systematic cognitive processes within human learning paradigms. However, the potential flaws of periodicity modeling in Transformer affect the learning efficiency and establishment of underlying principles from data for large language models (LLMs) built upon it. In this paper, we demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. We introduce FANformer, which integrates Fourier Analysis Network (FAN) into attention mechanism to achieve efficient periodicity modeling, by modifying the feature projection process of attention mechanism. Extensive experimental results on language modeling show that FANformer consistently outperforms Transformer when scaling up model size and training tokens, underscoring its superior learning efficiency. To further validate the effectiveness of FANformer, we pretrain a FANformer-1B on 1 trillion tokens. FANformer-1B exhibits marked improvements on downstream tasks compared to open-source LLMs with similar model parameters or training tokens. The results position FANformer as an effective and promising architecture for advancing LLMs. Our code and pretrained model are available at https://github.com/YihongDong/FANformer.
[…]
[ tikzsetting= fill=blueish, skipabove=0.33em, skipbelow=0.33em, linewidth=1pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=2pt, innerbottommargin=2pt, linecolor=gray95, roundcorner=2pt, shadowsize=4pt, shadowcolor=gray95 ]answerbox
FANformer: Improving Large Language Models Through Effective Periodicity Modeling
Yihong Dong1, Ge Li1,2, Xue Jiang1, Yongding Tao1, Kechi Zhang1, Hao Zhu1, Huanyu Liu1, Jiazheng Ding2, Jia Li ♂1, Jinliang Deng3, and Hong Mei1,4 1School of Computer Science, Peking University 2aiXcoder 3The Hong Kong University of Science and Technology 4Advanced Institute of Big Data dongyh@stu.pku.edu.cn, lige@pku.edu.cn
1 Introduction
In recent years, large language models (LLMs) have achieved remarkable progress across various natural language processing tasks, establishing themselves as a cornerstone of modern artificial intelligence Brown et al. (2020); Zhao et al. (2023); Minaee et al. (2024). The decoder-only Transformer architecture, in particular, has emerged as the de facto standard for LLM development due to its superior performance and scalability OpenAI (2023); DeepSeek-AI et al. (2024b); Groeneveld et al. (2024). Besides these advancements, Transformer-based models are also known for their immense demand for data and computational resources during training Kaplan et al. (2020); Hoffmann et al. (2022); Chowdhery et al. (2023). In comparison, humans are able to accomplish similar learning tasks with far fewer resources. This discrepancy suggests that existing LLM architectures still suffer from low learning efficiency, leaving substantial room for improvement in their ability to extract and generalize the knowledge from data.
Figure 1: The performance of FANformer on periodicity modeling and language modeling. (a) shows the training loss of Transformer and FANformer on the fitting mod function with scalar input and their performance at the 4,000th epoch. (b) shows the average performance of FANformer-1B and the open-source LLMs with comparable model parameters and training tokens.
Periodicity, characterized by recurring patterns, is a ubiquitous phenomenon in human life and learning processes Buzsaki (2006); Lake et al. (2017). The human brain leverages pattern recognition mechanisms to process information and acquire knowledge efficiently Zalta et al. (2020); Edalati et al. (2023); Zhan et al. (2018). However, general network architectures represented by Transformers have potential flaws in periodicity modeling, which could hinder their learning efficiency Dong et al. (2024b); Liu et al. (2020). As shown in Figure 1 (a), even for a simple mod function, Transformer demonstrates suboptimal performance despite being provided with sufficient training data and model capacity [^1] We sample 400K training data from the function of mod 5 and train a 110M parameter Transformer for 4K epochs.. This inefficiency can be exacerbated during the training process of LLMs to affect their performance, considering the periodicity hidden in large amounts of language data. Fourier Analysis Network (FAN) Dong et al. (2024b) has shown preliminary success in tasks with explicit or implicit periodic features, but its integration with Transformer architectures for large-scale language modeling remains an open challenge.
In this paper, we present FANformer, a novel foundation architecture for LLMs that integrates FAN into the attention mechanism of Transformer to improve the learning efficiency and performance of LLMs, by achieving efficient periodicity modeling. It leverages FAN to introduce Fourier principles for capturing and representing periodic patterns, thereby enhancing the Transformer’s capability to learn and generalize from data. Specifically, we modify the feature projection process of attention mechanism to incorporate frequency-domain representations to facilitate capturing and modeling periodicity. Figure 1 (a) demonstrates the significant advantages of FANformer over Transformer on periodicity modeling, with faster convergence speed and better results. In Figure 1 (b), we can observe that FANformer-1B achieves superior performance with higher utilization efficiency of model parameter and training token when benchmarked against comparable Transformer-based LLMs.
To comprehensively validate the effectiveness and scalability of FANformer, we conducted extensive experiments on language modeling tasks. The results of scaling both model parameters and training tokens highlight that FANformer consistently surpasses Transformer, requiring only 69.2% of model parameters or 79.7% of training tokens to achieve comparable performance. We also implement a complete pretraining pipeline to pretrain a 1.1-billion parameter FANformer (FANformer-1B) on 1 trillion tokens. Experiments on various downstream tasks demonstrate that FANformer-1B outperforms open-source LLMs of the same size with fewer training tokens, and exceeds LLMs with three times the parameters when using the same training token. Moreover, through further detailed analysis, we reveal that FANformer is a superior choice compared to other variant architectures and discover three interesting findings: 1) By observing the training process, we discover the notable enhancements in FANformer’s learning efficiency over Transformer as the model continues to learn from the data. 2) For larger FANformers, the optimal hyperparameter to control the proportion of periodicity modeling exhibits an increasing trend, suggesting that more powerful FANformers could extract richer hidden periodicity features. 3) FANformer facilitates the rule-based learning paradigm, effectively mitigating the occurrence of "holes" inherent in the case-based reasoning of Transformer Hu et al. (2024). These findings underscore the potential of FANformer as an effective and scalable architecture for advancing LLMs.
The main contributions of our work can be summarized as follows:
  • We first demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs.
  • We propose FANformer, a novel LLM architecture, which uses a simple yet effective approach to integrate FAN into attention mechanism for efficient periodicity modeling, consistently outperforming Transformers in scaling model parameters and training tokens.
  • We pretrain and open-source FANformer-1B, which surpasses state-of-the-art publicly available LLMs with similar parameter counts or training token budgets on downstream tasks.
Figure 2: Left: The illustration of FANformer’s architecture. Right: The pseudocode of Multi-head ATF, where p is the hyperparameter that controls the proportion of periodicity modeling for $\mathbf{X}_{p}$.
2 Preliminary Knowledge
Fourier Analysis Network (FAN) Dong et al. (2024b) enhances neural networks by introducing Fourier principles for effective periodicity modeling. The core component of FAN is its layer design, which combines periodic basis functions with standard linear transformations. Given a input $\mathbf{X}$, the FAN layer is defined as:
$\text{FANLayer}(\mathbf{X})=$ $[\cos(W_{p}\mathbf{X})\|\sin(W_{p}\mathbf{X})$
$\quad\|\sigma(W_{\bar{p}}\mathbf{X}+B_{\bar{p}})]$ (1)
[…]
3 FANformer
In this section, we provide a detailed description of FANformer for sequence modeling and adopt a decoder-only model to illustrate the architecture.
Given an input sequence $\boldsymbol{s}=\left\{s_{1},s_{2},\cdots,s_{l}\right\}\in\mathbb{R}^{l}$, where $s_{i}$ denotes the $i$-th token and $l$ represents the length of sequence $\boldsymbol{s}$, it is first mapped to the input embedding as $\mathbf{X}^{0}=[\mathbf{x}_{1},\mathbf{x}_{2},\cdots,\mathbf{x}_{l}]\in\mathbb{R}^{l\times d_{h}}$, where $d_{h}$ represents the model’s hidden dimension. The embedding is subsequently fed into the model to obtain the final output $\mathbf{X}^{N}$, with each $n$-th layer of FANformer processing $\mathbf{X}^{n-1}$, where $n\in[1,N]$. The core of each FANformer layer lies in a revised attention module that integrates FAN, referred to as the ATtention-Fourier (ATF) module.
3.1 ATF
The attention mechanism serves as a core component of Transformer architectures, enabling dynamic interaction between tokens through query-key-value (QKV) projections. While effective for general sequence modeling, its standard implementation exhibits limitations in capturing periodic patterns due to the inherent locality of linear projections in the time domain. To address this, we propose the ATF module, which incorporates the operations of FAN into the QKV projection process to explicitly model periodicity in the frequency domain. Specifically, given the input $\mathbf{X}\in\mathbb{R}^{l\times d_{h}}$, we first calculate $\mathbf{X}_{F}\in\mathbb{R}^{l\times d_{h}}$ as:
$\mathbf{X}_{F}$ $=\text{FANLayer}^{\prime}(\mathbf{X}),$
$=[\cos(W_{p}x)||\sin(W_{p}x)||(B_{\bar{p}}+W_{\bar{p}}x)],$ (2)
where $\text{FANLayer}^{\prime}$ represents a variant of Eq. (1) with the activation function $\sigma$ in Eq. (1) replaced by the identity function, i.e., $\sigma(x)=x$, in this paper. On this basis, we employ the linear transform to $\mathbf{X}_{F}$ to compute QKV projections, i.e., $\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}\in\mathbb{R}^{l\times d_{h}}$, as follows:
$[\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}]=\mathbf{X}_{F}[\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}],$ (3)
where $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{d_{h}\times d_{h}}$ are learnable parameters. Similar to the standard attention mechanism, the computation of ATF is defined as:
$\text{ATF}(\mathbf{X})=\text{softmax}\left(\frac{\mathbf{Q}_{F}\mathbf{K}_{F}^{\top}}{\sqrt{d_{h}}}\right)\mathbf{V}_{F},$ (4)
where $\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}$ are computed using the input $\mathbf{X}$ via Eq. (2) and Eq. (3). To enhance the model’s capacity, we extend the ATF module to multiple heads. Given input $\mathbf{X}\in\mathbb{R}^{l\times d_{h}}$, the Multi-head ATF first projects $\mathbf{X}$ into $k$ independent heads through the ATF module. For the $i$-th head, we have:
$\text{Head}_{i}=\text{ATF}(\mathbf{X}|\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i},\mathbf{W}_{V}^{i};k),$ (5)
where $\mathbf{W}_{Q}^{i},\mathbf{W}_{K}^{i},\mathbf{W}_{V}^{i}\in\mathbb{R}^{d_{h}\times d_{k}}$ are learnable parameters for query, key, and value projections respectively, with $d_{k}=d_{h}/k$. The outputs of all heads are concatenated and linearly transformed:
$\text{MultiHeadATF}(\mathbf{X})=[\text{Head}_{1}\|...\|\text{Head}_{k}]\mathbf{W}_{O},$ (6)
where $\mathbf{W}_{O}\in\mathbb{R}^{d_{h}\times d_{h}}$ is the learnable parameter of out projection matrix. Note that $\text{ATF}(\mathbf{X})$ is mathematically equivalent to $\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$ (the detailed derivations are provided in Appendix F). This equivalence enables a simple yet effective implementation of Multi-head ATF as shown in Figure 2, which can seamlessly incorporate various advancements in traditional attention mechanisms, such as FlashAttention Dao et al. (2022).
3.2 Overall Architecture
The FANformer model comprises $N$ stacked FANformer layers, where each FANformer layer consists of a Multi-head ATF module and a feed-forward network (FFN) module. Following the previous work Groeneveld et al. (2024); Touvron et al. (2023b), we employ SwiGLU Ramachandran et al. (2018); Shazeer (2020) and pre-norm Zhang et al. (2019) as the enhancements to Transformer-based LLMs. Specifically, the $n$-th FANformer layer can be defined as:
$\mathbf{Y}^{n}=\text{MultiHeadATF}(\text{Norm}(\mathbf{X}^{n}))+\mathbf{X}^{n},$ (7)
$\mathbf{X}^{n+1}=\text{FFN}(\text{Norm}(\mathbf{Y}^{n}))+\mathbf{Y}^{n},$ (8)
where the MultiHeadATF module is computed via Eq. (6) and the FFN module, which leverages the SwiGLU activation, is expressed as:
$\text{FFN}(\mathbf{X})=(\text{Swish}(\mathbf{X}\mathbf{W}_{1})\otimes\mathbf{X}\mathbf{W}_{2})\mathbf{W}_{3},$ (9)
where $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d_{h}\times d_{f}}$, $\mathbf{W}_{3}\in\mathbb{R}^{d_{f}\times d_{h}}$ are learnable parameters, $\otimes$ denotes element-wise multiplication, and $d_{f}$ is the intermediate dimension. The overview of FANformer’s architecture is illustrated in Figure 2.
4 Evaluation
We begin with the implementation details of our experiments (Section 4.1), followed by a comprehensive evaluation of FANformer from three distinct perspectives: First, we investigate the scalability of FANformer by examining its performance trends on language modeling tasks with respect to model size and training tokens (Section 4.2). Second, we evaluate the capabilities of the pre-trained FANformer-1B model across multiple downstream tasks (Section 4.3). Third, we conduct an in-depth empirical analysis of FANformer, including ablation study, learning efficiency, hyperparameter impact, reasoning mechanism, etc. (Section 4.4). See Appendix A-E for more experiments.
4.1 Implementation Details
The experiments in this paper are conducted on eight GPUs of Tesla A100-PCIe-40G. We adopt the open language model OLMo Groeneveld et al. (2024) as the baseline Transformer. Building upon this foundation, we apply ATF into OLMo to construct our FANformer, where hyperparameter p is set to 0.25 by default. For pertaining FANformer-1B, we sample 1T training tokens from OLMo’s training data, i.e., Dolma Soldaini et al. (2024). For other experiments, we train models on a smaller sample of Dolma, i.e., Dolma v1_6-sample (AllenAI, 2023), with roughly 10B tokens. The detailed pretraining and experimental setups are provided in Appendix G.
Table 1: Zero-shot performance of FANformer-1B versus other comparable open-source LLMs on 8 core tasks from the downstream evaluation suite following OLMo Groeneveld et al. (2024). The results of baselines are taken from the work Groeneveld et al. (2024); Ye et al. (2024); Dong et al. (2024a).
Models Param. Tokens ARC-C ARC-E BoolQ Hella. OBQA PIQA SCIQ Wino. Avg.
LLMs around 1B parameters
Qwen2.5-1.5B 1.5B 18T 41.2 75.5 74.0 50.2 52.4 75.7 94.7 63.3 65.9
R1-Distill-Qwen1.5B 1.5B 18T+ 36.2 54.4 69.1 41.8 35.4 65.1 89.5 55.3 55.9
Llama-3.2-1B 1.2B 9T 31.4 65.6 64.3 47.8 46 74.5 92.3 60.7 60.4
TinyLlama-v1.1-3T 1.1B 3T 34.8 53.2 64.6 58.7 43.6 71.1 90.5 58.9 59.4
OLMo-1B 1.1B 2T 34.5 58.1 60.7 62.5 46.4 73.7 88.1 58.9 60.4
[…]
OpenLLaMA-v2-3B 3B 1T 33.9 67.6 65.7 70.0 26 76.7 92.9 62.9 62.0
StableLM-base-alpha-v2-3B 3B 1T 32.4 67.3 64.6 68.6 26.4 76 89.5 62.1 60.9
TinyLlama-v1.1-1T 1.1B 1T 33.1 49.5 58.4 52.5 37.8 70.4 86.4 55.2 55.4
[…]
4.2 Scalability of FANformer
We explore the scalability of FANformer compared with Transformer to investigate performance trends in the construction of much larger models.
Setup.
We follow OLMo’s configuration and vary the FFN’s intermediate dimension $d_{f}$ to keep the number of parameters consistent for all models in this experiment. For scaling up mode parameters, we adopt Dolma v1_6-sample as training data and train LLMs from 268M to 7B. We compare FANformer with the standard Transformer and a variant of FANformer, termed Transformer+ATM, which uses MLP layer instead of FAN layer in FANformer. For scaling up training tokens, we train 1B LLMs on the first 200 billion of our sampled 1T tokens.
Figure 3: Language modeling loss of scaling up model parameter and training tokens. Left: we train LLMs from 268M to 7B parameters. Right: we evaluate 1.0B LLMs every 20B tokens up to 200B tokens.
Results.
As shown in Figure 3, the scaling law (Kaplan et al., 2020) empirically aligns well with the results obtained from our FANformer, underscoring its superior scalability properties. Figure 3 (left) reveals that the implementation of FAN consistently surpasses the performance of the standard Transformer across a range of model sizes. This finding highlights FAN’s enhanced scalability in terms of parameter efficiency, as it achieves comparable performance with only 69.2% of the parameters required by the standard Transformer. Notably, the scaling curve of Transformer+ATM closely overlaps with that of the standard Transformer, indicating that merely revising attention mechanisms using MLP Layer is insufficient. This observation further emphasizes that FANformer’s performance gains are not attributable to network depth increase, but rather to its special architectural design. Figure 3 (right) demonstrates that FANformer achieves performance parity with the standard Transformer while utilizing significantly fewer training tokens. Specifically, FANformer requires only 159.6B training tokens to match the performance of the standard Transformer trained on 200B tokens, representing a 20.3% reduction in training resource requirements. These findings suggest that FANformer exhibits superior utilization efficiency in terms of both model parameters and training tokens compared to the standard Transformer architecture.
4.3 Performance of FANformer-1B
We pretrain FANformer-1B on 1 trillion tokens and report zero-shot performance on a set of commonsense downstream tasks, following previous work (Brown et al., 2020; Touvron et al., 2023a; Groeneveld et al., 2024, inter alia).
Setup.
The downstream evaluation suite consists of 8 core tasks, including ARC-C (Clark et al., 2018), ARC-E (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA Mihaylov et al. (2018), PIQA (Bisk et al., 2020), SCIQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2020). We compare pretrained FANformer-1B to seven open-source LLMs with comparable model parameters or training tokens, including Qwen2.5-1.5B (Team, 2024), R1-Distill-Qwen1.5B DeepSeek-AI et al. (2025), Llama-3.2-1B Dubey et al. (2024), TinyLlama-v1.1-1B Zhang et al. (2024), OLMo-1B (Groeneveld et al., 2024), OpenLLaMA-v2-3B Geng et al. (2023), and StableLM-base-alpha-v2-3B Tow (2023).
Table 2: Results of ablation study and variant analysis on LLMs with 1B parameters trained on Dolma v1_6-sample dataset (about 10B tokens). The complete experimental results can be found in Table 4 and Table 5 of Appendix.
Variants Param. Training Loss $\downarrow$ V2 Eval Loss $\downarrow$ V2 Eval PPL $\downarrow$ V3 Eval Loss $\downarrow$ V3 Eval PPL $\downarrow$ DownStream Avg Acc. $\uparrow$
Transformer 1.0×1.0$\times$1.0 × 2.889 3.33 30.20 3.07 24.28 53.10
Transformer (FFN $\leftarrow$ FAN) 1.0×1.0$\times$1.0 × 2.880 3.31 29.79 3.05 23.96 53.95
[…]
Transformer + ATM 1.0×1.0$\times$1.0 × 2.890 3.33 30.31 3.07 24.36 53.69
Transformer + ATL 1.0×1.0$\times$1.0 × 2.882 3.31 29.68 3.05 23.97 53.46
FANformer + Activation 1.0×1.0$\times$1.0 × 2.893 3.34 30.64 3.07 24.50 53.61
FANformer 1.0×1.0$\times$1.0 × 2.863 3.30 29.40 3.04 23.62 55.19
[…]
Transformer + ATM 1.06×1.06$\times$1.06 × 2.886 3.33 30.18 3.06 24.28 52.86
Transformer + ATL 1.06×1.06$\times$1.06 × 2.879 3.31 29.76 3.05 23.94 54.23
FANformer + Activation 1.04×1.04$\times$1.04 × 2.887 3.34 30.57 3.07 24.39 53.13
FANformer 1.04×1.04$\times$1.04 × 2.856 3.29 29.22 3.03 23.47 54.88
Results.
Table 1 presents the evaluation results of our pre-trained FANformer-1B on downstream tasks. It is evident that FANformer-1B surpasses LLMs with comparable parameter sizes, such as Llama-3.2-1B, TinyLlama-v1.1-3T, and OLMo-1B, while utilizing significantly fewer training data. Compared to the base model OLMo-1B, FANformer-1B achieves a relative improvement of 8.8% in the average performance of downstream tasks using only half the training data. On these tasks, FANformer-1B also demonstrates performance comparable to Qwen2.5-1.5B, which is the current SOTA LLM around 1B. For LLMs training on 1T tokens, FANformer-1B even exceeds LLMs with three times the parameters, showing an average relative performance improvement of 6.0-7.9% across all tasks. Moreover, while R1-Distill-Qwen1.5B shows notable improvements in reasoning capabilities based on its reported performance, it exhibits significantly weaker general performance on these commonsense downstream tasks. This observation shows the shortcomings of distillation, highlighting the necessity of the pre-training stage and the importance of research into more efficient model architectures.
4.4 Further Analysis
4.4.1 Ablation Study and Variant Analysis
Setup.
We compare FANformer to other variant architectures, including the above-mentioned Transformer+ATM, Transformer+ATL: use two linear transforms to compute QKV projection, FANformer +Activation: employ Eq. (1) with GELU Hendrycks et al. (2016) activation function instead of Eq. (2), Transformer (FFN $\leftarrow$ FAN): replace FFN with FAN Dong et al. (2024b), and standard Transformer as their ablations.
Results.
From Table 4.3, we have the following findings: 1) FANformer consistently outperforms other variant architectures in both scenarios of the same parameter and same dimension on all evaluation metrics. 2) The performance of Transformer+ATM and Transformer+ATL is notably inferior to that of FANformer, indicating that the core improvement stems from the ATF module we designed. 3) Although Transformer (FFN $\leftarrow$ FAN) yields some improvement, this enhancement is inferior to the gains achieved by FANformer, suggesting that integrating periodicity modeling within attention is more advantageous than FFN on language modeling. 4) Incorporating activation functions such as GELU into the attention mechanism tends to degrade model performance. Specifically, FANformer+Activation and Transformer+ATM exhibit weaker performance compared to FANformer and Transformer+ATL, likely because these activation functions suppress certain features, thereby hindering subsequent attention operations.
Figure 4: Training loss of FANformer and Transformer on early training steps. The complete training loss is provided in Figure 8 of Appendix.
4.4.2 Training Dynamics
We perform a comparative analysis of the loss trends during the training process between our FANformer and Transformer, as illustrated in Figure 4. The experimental results indicate that the loss of FANformer decreases more slowly in the early stages compared to Transformer, which we hypothesize may be due to the initial lack of established periodic modeling. As the training progresses and periodic modeling gradually improves, FANformer demonstrates a faster convergence rate, with its loss decreasing more rapidly than that of Transformer. This result suggests that as the model progressively learns from the data, the learning efficiency of our FANformer notably surpasses the standard Transformer.
(a) Modular Addition Task
(b) Linear Regression Task
Figure 5: Performance of FANformer and Transformer on modular addition and linear regression tasks, where the darkened regions indicate areas where the model performance approaches zero, signifying the emergence of the "hole" as described in the work Hu et al. (2024).
Figure 6: Effect of hyperparameter $p$ in FANformer on its training loss and downstream task performance, where the red dashed line represents the training loss of Transformer, while the blue dashed line denotes the performance on downstream tasks of Transformer.
4.4.3 Effect of hyperparameter $p$
We systematically investigate the impact of hyperparameter p, which controls the proportion of periodicity modeling in FANformer, on model performance across its value range. The experimental results from the 1B-scale FANformer (as shown in Figure 6) demonstrate that our model exhibits strong robustness in terms of training loss and downstream task accuracy, with relatively small performance fluctuations. Furthermore, regardless of the variation in p values, FANformer consistently outperforms the standard Transformer (horizontal baseline). Analysis of experimental results from models of different scales (300M, 1B, 3B) (as illustrated in Figure 9 of Appendix) reveals a clear trend: larger models tend to exhibit higher optimal p values. This observation suggests that more powerful FANformers are better equipped to extract more intricate latent periodicity features.
4.4.4 Case-based and Rule-based Reasoning
Setup.
Following the work Hu et al. (2024), we evaluate the case-based and rule-based reasoning of Transformer and our FANformer on two tasks, including: (1) Modular addition: $c=(a+b)\mod 113$ with $a,b\in[0,112]$; (2) Linear regression: $c=a+2b+3$ with $a,b\in[0,99]$. We finetune pretained LLMs, i.e., OLMo-1B and FANformer-1B, on each task dataset for 500 epochs and their performance is measured via the Leave-Square-Out method, sampling 10 generations at temperature 0.5 per test point.
Figure 7: Training accuracy and test accuracy of FANformer and Transformer on modular addition and linear regression tasks.
Results.
As illustrated in Figure 5, both Tranformer and our FANformer achieve near-perfect accuracy on the training set of modular addition and linear regression tasks, approaching approximately 100%. However, a critical divergence emerges in their performance on the test sets. Specifically, as shown in Figure 5, Transformer exhibits a pronounced failure to generalize, resulting in a "black hole" in the center of the figure, indicating that its accuracy on the test dataset drops to nearly zero. This observation is consistent with the findings reported in the work  (Hu et al., 2024). In contrast, FANformer demonstrates a marked improvement in addressing the "hole" issue. In the linear regression and modular addition tasks, there is no obvious hole observed, further corroborating the hypothesis that, relative to the Transformer-based model, FANformer possesses a stronger tendency to learn underlying rules, thereby achieving superior generalization performance.
5 Related Work
In this section, we outline the three most relevant directions and associated papers of this work.
5.1 Large Language Models
The rapid advancement of language models has revolutionized natural language processing and artificial intelligence research Radford (2018); Dubey et al. (2024); DeepSeek-AI et al. (2025). The emergence of GPT-3 Brown et al. (2020), with 175B parameters, showcased remarkable few-shot prompting abilities, suggesting that scaling laws Kaplan et al. (2020) could unlock emergent capabilities. Recent notable LLMs like PaLM Chowdhery et al. (2023), LLaMA Touvron et al. (2023a), GPT-4 Chowdhery et al. (2023), and DeepSeek Bi et al. (2024) further pushed the boundaries of model size and performance. Additionally, the open-source release of OLMo Groeneveld et al. (2024) has provided valuable resources for the community, enabling more accessible training of LLMs.
5.2 Advances in Transformer Architecture
Recent advancements in Transformer architecture primarily address two fundamental limitations: computational inefficiency in long-context processing and insufficient expressiveness of attention mechanisms. To tackle long-context processing limitations, Sparse Transformer (Child et al., 2019) and Longformer (Beltagy et al., 2020) introduce sparsity and local Attention. Query mechanism innovations like MQA (Shazeer, 2019) and GQA Ainslie et al. (2023) optimize token interaction efficiency. For inference acceleration, MLA DeepSeek-AI et al. (2024a) pioneers a low-rank latent space compression technique for Key-Value pairs, dramatically reducing cache memory requirements. Hardware-level optimizations emerge through FlashAttention Dao et al. (2022), which strategically minimizes GPU memory access overhead during attention computation. To improve the expressiveness of networks, Probabilistic Attention Keys Nguyen et al. (2022) replace deterministic key-query interactions with Gaussian kernels, enabling better capture of semantic relationships. Selective Attention Leviathan et al. (2024) introduces dynamic token masking to suppress irrelevant features to refine attention mechanism. Differential Transformer Ye et al. (2024) addresses attention noise in long contexts by computing dual softmax maps and performing subtraction.
Different from previous work, we improve language modeling by addressing the challenge of modeling periodicity in Transformers, which can seamlessly incorporate the aforementioned works for revising the attention mechanism, as demonstrated in the derivation provided in Appendix F.
5.3 Fourier-based Neural Networks
Previous research on Fourier-based Neural Networks was aimed at solving some domain-specific applications Zuo et al. (2005); Tan (2006); Chen et al. (2022); Li et al. (2021). Some studies specifically explored the use of sinusoidal activations (e.g., cosine Silvescu (1999) Ngom et al. (2021) or sine Parascandolo et al. (2016); Sitzmann et al. (2020)) to approximate periodic patterns Liu (2013). However, these approaches lacked generalizability beyond narrow domains due to rigid frequency parameterization and limited scalability Uteuliyeva et al. (2020); Liu et al. (2020). FAN (Dong et al., 2024b) introduces Fourier Principle into neural networks to effectively solve periodicity modeling challenges while maintaining broad applicability similar to MLP. FNet Lee-Thorp et al. (2022) replaces self-attention with Fourier Transform to achieve linear complexity, but it sacrifices the performance of language modeling. In contrast, we employ effective periodicity modeling to improve LLMs.
6 Conclusion
We propose FANformer, a novel LLM architecture that enhances learning efficiency by integrating Fourier Analysis Network into self-attention mechanism for effective periodicity modeling. Experiments demonstrate that FANformer outperforms Transformer when scaling model parameters and training tokens, achieving better performance with 31% fewer parameters and 20% fewer tokens. Pretrained FANformer-1B surpasses open-source LLMs of comparable size or training scale on various downstream tasks, highlighting the benefits of periodicity-aware architecture design. The discovery of FANformer’s enhanced scalability, learning efficiency, and rule-based learning advantages suggests potential pathways for developing more efficient and high-performance language models.
7 Limitations
Our work has several limitations, which we aim to address in our future work:
First, due to constraints in computational resources, we only pretrain the FANformer-1B on 1 trillion tokens. However, our experimental results regarding FANformer’s scalability indicate that our FANformer demonstrates favorable scaling behavior during training, suggesting that increasing the model size and training tokens could lead to more significant performance improvements. To explore this further, we plan to seek additional computational resources to train larger-scale language models.
Second, our work is orthogonal to the existing approaches for revising the attention mechanism, i.e., our work can seamlessly incorporate them, as verified in the derivation provided in Appendix F. There are numerous variants of attention mechanisms, as discussed in the related work (Section 5.2), such as Flash Attention Dao et al. (2022), MQA Shazeer (2019), and MLA DeepSeek-AI et al. (2024a). In this work, we only incorporate Flash Attention for necessary acceleration, while leaving the exploration of other approaches for future work.
Third, although we have observed that enhancing the ability of language models to model periodic patterns can improve language modeling performance, the underlying mechanisms responsible for this improvement remain underexplored. To the best of our knowledge, it has hardly been studied the role of periodicity or the potential periodic behaviors of LLMs on language modeling. Therefore, in future work, we will conduct a more comprehensive investigation into the fundamental mechanisms of periodicity in language modeling.
Appendix A Training Loss Curves of OLMO and FANformer
We present the training loss curves for OLMO and FANformer trained on 1 trillion tokens (i.e., 250K steps) in Figure 8.
(a) Training loss of OLMO-1B
(b) Training loss of FANformer-1B
Figure 8: The training process of OLMO and FANformer. The data in Figure (a) is sourced from the publicly available results of OLMO (https://wandb.ai/ai2-llm/OLMo-1B?nw=nwuserdirkgr).
Appendix B Extended Results about Hyperparameter $p$ for Section 4.4.3
We investigate the impact of hyperparameter p on training loss across FANformer models of varying scales, with the results illustrated in Figure 9.
Figure 9: The impact of hyperparameter $p$ on FANformer models of varying sizes.
Appendix C Instruction Following with SFT
Model MMLU AlpacaEval ToxiGen TruthfulQA (Info+True)
0-shot $\uparrow$ %win $\uparrow$ % Toxic $\downarrow$ Accuracy $\uparrow$
OLMo-1B-SFT 24.3 1.90 2.8 55.8
FANformer-1B-SFT 26.7 2.51 10.4 83.0
Table 3: Evaluation results of OLMo-1B-SFT and FANformer-1B-SFT on MMLU, AlpacaEval, ToxiGen, and TruthfulQA (Info+True). Higher values are better for MMLU, AlpacaEval, and TruthfulQA, while lower values are better for ToxiGen.
C.1 Models
FANformer-1B-SFT: Our pretrained model on 1 trillion tokens, fine-tuned using supervised fine-tuning (SFT) on the tulu-3-sft-olmo-2-mixture dataset.
OLMo-1B-SFT: A 1B parameter version of OLMo, pre-trained on 3 trillion tokens and fine-tuned using supervised fine-tuning (SFT) on the tulu-3-sft-olmo-2-mixture dataset. Model available at allenai/OLMo-1B-hf.
For both models, we follow the tokenizer’s chat template for prompt formatting when available.
C.2 Evaluation Setup
We evaluate the models on four benchmarks: MMLU, TruthfulQA, AlpacaEval, and ToxiGen. The evaluation is conducted using the open-instruct evaluation suite, which has been widely adopted for evaluating instruction-tuned language models. Below, we describe the setup for each benchmark.
MMLU
We use the official MMLU evaluation script with 0-shot prompting. The maximum token length is set to 2048, and we do not employ chain-of-thought (CoT) reasoning. The evaluation reports the average accuracy across test examples.
AlpacaEval
We use the AlpacaEval V1 benchmark with the default setup from the official repository  [^2] https://github.com/tatsu-lab/alpaca_eval. The evaluated models generate responses for 805 prompts, and GPT-4 is employed to compare the responses with those from the reference model (gpt-4-1106-preview). Given the updates in the AlpacaEval repository, the default models have changed over time. Currently, the default setup uses the weighted_alpaca_eval_gpt4_turbo annotator as the annotator and gpt-4-1106-preview as the reference model. Therefore, our evaluation aligns with the current default configuration in the official AlpacaEval repository.
ToxiGen
For ToxiGen, we focus on the prompts specifically designed to elicit toxic language (‘hateful’ prompts). To reduce evaluation costs, we use a subset of 500 prompts per group for testing. The toxicity classifier used is toxigen_roberta. We report the percentage of generations classified as toxic by the classifier.
TruthfulQA
For TruthfulQA, we use the generation setting with the default QA prompt format, including 6 in-context examples. The judge model for evaluating truthfulness and informativeness is allenai/truthfulqa-truth-judge-llama2-7B, which is adopted in the open-instruct evaluation suite and OLMo’s evaluation. We report the percentage of responses deemed both informative and truthful.
Appendix D Detailed Results of Ablation Study for Section 4.4.1
In ablation study, we report the average results across various tasks on V2 Validation Sets, V3 Validation Sets, and Downstream tasks, with the specific tasks detailed in Section G.3. The complete results are detailed in Table 4 and Table 5.
Table 4: The detailed results of ablation study (Part One). All models keep the same number of parameters and are pretrained on Dolma v1_6-sample dataset (about 10B tokens).
Transformer Transformer + ATM Transformer + ATL FANformer + Activation FANformer
V2 Validation Sets 4chan Loss 2.68 2.68 2.66 2.70 2.66
PPL 14.60 14.53 14.36 14.88 14.34
c4_100_domains Loss 3.11 3.11 3.10 3.12 3.08
PPL 22.38 22.52 22.18 22.63 21.87
c4_en Loss 3.27 3.28 3.27 3.29 3.25
PPL 26.40 26.54 26.22 26.78 25.85
gab Loss 3.90 3.90 3.89 3.91 3.87
PPL 49.58 49.64 49.11 50.05 47.83
ice Loss 3.20 3.21 3.19 3.21 3.17
PPL 24.59 24.77 24.25 24.82 23.93
m2d2_s2orc Loss 3.56 3.57 3.56 3.59 3.56
PPL 35.34 35.47 34.99 36.24 35.05
m2d2_wiki Loss 3.14 3.14 3.13 3.15 3.11
PPL 23.17 23.13 22.90 23.29 22.48
manosphere Loss 3.47 3.48 3.46 3.48 3.45
PPL 32.21 32.46 31.85 32.62 31.48
mc4_en Loss 3.02 3.02 3.01 3.03 2.99
PPL 20.53 20.52 20.22 20.76 19.91
pile Loss 2.76 2.76 2.74 2.77 2.73
PPL 15.84 15.74 15.53 15.99 15.30
ptb Loss 3.68 3.70 3.64 3.71 3.66
PPL 39.68 40.51 38.23 40.74 38.75
twitterAEE Loss 4.10 4.10 4.07 4.11 4.07
PPL 60.25 60.18 58.79 61.10 58.54
wikitext_103 Loss 3.33 3.33 3.30 3.35 3.29
PPL 28.03 28.07 27.15 28.48 26.88
Average Loss 3.33 3.33 3.31 3.34 3.30
PPL 30.20 30.31 29.68 30.64 29.40
[…]
PPL 24.80 24.86 24.60 25.04 24.24
dolma_books Loss 3.56 3.56 3.54 3.57 3.52
PPL 34.98 35.32 34.43 35.57 33.96
dolma_common-crawl Loss 3.23 3.24 3.23 3.24 3.21
PPL 25.32 25.42 25.16 25.47 24.76
dolma_pes2o Loss 2.86 2.85 2.84 2.86 2.83
PPL 17.45 17.35 17.09 17.53 16.88
dolma_reddit Loss 3.44 3.44 3.43 3.45 3.42
PPL 31.13 31.35 30.94 31.42 30.54
dolma_stack Loss 1.42 1.41 1.40 1.42 1.39
PPL 4.13 4.10 4.06 4.13 4.01
dolma_wiki Loss 3.04 3.04 3.03 3.04 3.01
PPL 20.89 20.84 20.62 20.97 20.26
ice Loss 3.19 3.20 3.18 3.20 3.17
PPL 24.41 24.56 24.09 24.63 23.75
m2d2_s2orc Loss 3.70 3.70 3.69 3.70 3.68
PPL 40.35 40.61 40.22 40.56 39.50
pile Loss 2.74 2.73 2.72 2.75 2.70
PPL 15.44 15.35 15.16 15.58 14.92
wikitext_103 Loss 3.34 3.34 3.31 3.35 3.30
PPL 28.21 28.21 27.33 28.57 27.03
Average Loss 3.07 3.07 3.05 3.07 3.04
PPL 24.28 24.36 23.97 24.50 23.62
[…]
hellaswag ACC 33.87 33.84 34.02 33.75 34.37
winogrande ACC 52.80 51.62 49.96 48.78 51.72
openbook_qa ACC 28.00 28.20 28.00 28.20 29.00
sciq ACC 70.30 72.10 69.00 67.20 71.80
arc_easy ACC 45.44 46.14 47.19 47.02 45.61
copa ACC 62.00 66.00 65.00 66.00 66.00
rte ACC 51.26 52.35 52.71 48.74 57.04
commitment_bank ACC 42.86 41.07 46.43 53.57 44.64
mrpc ACC 81.05 81.22 81.22 81.22 81.47
sst2 ACC 50.11 51.49 49.08 49.08 59.11
Average ACC 53.10 53.69 53.46 53.61 55.19
Table 5: The detailed results of ablation study (Part Two). All models keep the same dimension and are pretrained on Dolma v1_6-sample dataset (about 10B tokens).
Transformer Transformer + ATM Transformer + ATL FANformer + Activation FANformer
V2 Validation Sets 4chan Loss 2.68 2.68 2.67 2.68 2.66
PPL 14.60 14.54 14.43 14.63 14.29
c4_100_domains Loss 3.11 3.11 3.10 3.11 3.08
PPL 22.38 22.43 22.11 22.49 21.69
c4_en Loss 3.27 3.28 3.26 3.28 3.24
PPL 26.40 26.51 26.12 26.54 25.61
gab Loss 3.90 3.90 3.89 3.91 3.87
PPL 49.58 49.41 48.97 50.11 47.79
ice Loss 3.20 3.20 3.19 3.21 3.17
PPL 24.59 24.62 24.22 24.90 23.69
m2d2_s2orc Loss 3.56 3.58 3.56 3.58 3.54
PPL 35.34 35.73 35.17 35.78 34.58
m2d2_wiki Loss 3.14 3.14 3.13 3.14 3.10
PPL 23.17 23.04 22.81 23.10 22.27
manosphere Loss 3.47 3.48 3.46 3.48 3.45
PPL 32.21 32.44 31.78 32.43 31.36
mc4_en Loss 3.02 3.02 3.01 3.03 2.99
PPL 20.53 20.51 20.22 20.61 19.86
pile Loss 2.76 2.76 2.74 2.77 2.72
PPL 15.84 15.78 15.54 15.90 15.24
ptb Loss 3.68 3.67 3.67 3.73 3.63
PPL 39.68 39.19 39.15 41.67 37.82
twitterAEE Loss 4.10 4.10 4.08 4.11 4.07
PPL 60.25 60.19 59.12 60.97 58.62
wikitext_103 Loss 3.33 3.33 3.31 3.34 3.29
PPL 28.03 27.96 27.29 28.22 26.98
Average Loss 3.33 3.33 3.31 3.34 3.29
PPL 30.20 30.18 29.76 30.57 29.22
[…]
PPL 24.80 24.78 24.52 24.82 24.00
dolma_books Loss 3.56 3.56 3.54 3.56 3.52
PPL 34.98 35.10 34.41 35.24 33.64
dolma_common-crawl Loss 3.23 3.23 3.22 3.23 3.20
PPL 25.32 25.25 25.09 25.35 24.55
dolma_pes2o Loss 2.86 2.85 2.84 2.86 2.82
PPL 17.45 17.37 17.12 17.44 16.79
dolma_reddit Loss 3.44 3.44 3.43 3.44 3.41
PPL 31.13 31.22 30.83 31.28 30.31
dolma_stack Loss 1.42 1.41 1.40 1.42 1.39
PPL 4.13 4.09 4.07 4.13 4.02
dolma_wiki Loss 3.04 3.03 3.03 3.04 3.00
PPL 20.89 20.78 20.61 20.88 20.10
ice Loss 3.19 3.20 3.18 3.21 3.16
PPL 24.41 24.44 24.04 24.72 23.55
m2d2_s2orc Loss 3.70 3.70 3.69 3.70 3.67
PPL 40.35 40.50 39.99 40.56 39.17
pile Loss 2.74 2.73 2.72 2.74 2.70
PPL 15.44 15.39 15.17 15.50 14.87
wikitext_103 Loss 3.34 3.34 3.31 3.35 3.30
PPL 28.21 28.12 27.46 28.36 27.12
Average Loss 3.07 3.06 3.05 3.07 3.03
PPL 24.28 24.28 23.94 24.39 23.47
[…]
hellaswag ACC 33.87 33.96 34.22 33.92 35.15
winogrande ACC 52.80 51.62 50.12 51.07 51.38
openbook_qa ACC 28.00 28.00 28.80 28.60 28.40
sciq ACC 70.30 70.90 70.40 70.20 70.30
arc_easy ACC 45.44 48.60 47.02 44.91 48.95
copa ACC 62.00 67.00 67.00 65.00 69.00
rte ACC 51.26 51.99 54.87 54.51 54.87
commitment_bank ACC 42.86 32.14 41.07 37.50 39.29
mrpc ACC 81.05 81.17 80.59 81.17 81.11
sst2 ACC 50.11 50.92 55.73 51.15 60.55
Average ACC 53.10 52.86 54.23 53.13 54.88
Appendix E Extended results of Section 4.4.4
The training and testing performance metrics, including loss and accuracy, for case-based and rule-based reasoning are presented in Figure 10 and Figure 11, respectively.
(a) Training loss on modular addition task
(b) Training loss of linear regression task
Figure 10: Training loss of FAN and Transformer on case-based and rule-based reasoning.
(a) Training accuracy on modular addition task
(b) Test accuracy of modular addition task
(c) Training accuracy on linear regression task
(d) Test accuracy of linear regression task
Figure 11: Accuracy of FAN and Transformer during training and testing on case-based and rule-based reasoning.
Appendix F Derivation of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
[…]
$=\text{Attention}(\mathbf{X}_{F})\quad\text{{(matches standard attention: $\text{Attention}(\mathbf{Z})=\text{softmax}\left(\frac{\mathbf{Z}\mathbf{W}_{Q}(\mathbf{Z}\mathbf{W}_{K})^{\top}}{\sqrt{d_{k}}}\right)\mathbf{Z}\mathbf{W}_{V}$)}}$
[…]
Appendix G Comprehensive Experimental Details
G.1 Detailed training settings of FANformer
We train FANformer-1B using the ZeRO optimizer strategy (Rajbhandari et al., 2020) via PyTorch’s DDP framework (Li, 2018). Following OLMo Groeneveld et al. (2024), we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens). To improve throughput, we employ PyTorch’s amp module with the bfloat16 format. We employ the AdamW optimizer Loshchilov et al. (2019) for the model’s training process. The learning rate for all LLMs is set to 4.0e-4. We warm up the learning rate over 2000 steps ( 8B tokens) and then decay it in a cosine manner from there down to a tenth of the peak learning rate over the remainder of training. We employ FlashAttention Dao et al. (2022) to accelerate the model training and inference processes, leveraging its ability to optimize memory usage and computational efficiency. The total GPU computational cost for pre-training FANformer-1B amounts to approximately 4,700 GPU hours.
G.2 Detailed Setup for Section 4.2
For different model sizes in Figure 3, the hidden dimension, number of layers, and number of heads are listed in Table 6.
Table 6: Model size and setup used for FANformer in Section 4.2, where Transformers follows the setups of previous work OLMo Groeneveld et al. (2024).
Model Size Hidden Dim. Num Layers Num Heads Weight Tying
FANformer-300M 268M 1024 16 16 True
FANformer-600M 604M 1536 16 16 True
FANformer-1B 1.1B 2048 16 16 True
FANformer-3B 2.6B 2560 24 20 False
FANformer-7B 6.7B 4096 24 32 False
G.3 Validation Set And Downstream Tasks
Following (Zhu et al., 2024), we use V2 Validation Sets, V3 Validation Sets, and Downstream tasks to evaluate our approach. The specific tasks included in V2 validation sets, V3 validation sets, and downstream tasks are listed in Table 7.
Table 7: Validation Set And Downstream Tasks.
V2 Validation Sets
v2-small-4chan-validation
v2-small-c4_100_domains-validation
v2-small-c4_en-validation
v2-small-gab-validation
v2-small-ice-validation
v2-small-m2d2_s2orc-validation
v2-small-m2d2_wiki-validation
v2-small-manosphere-validation
v2-small-mc4_en-validation
v2-small-pile-validation
v2-small-ptb-validation
v2-small-twitterAEE-validation
v2-small-wikitext_103-validation
V3 Validation Sets
v3-small-c4_en-validation
v3-small-dolma_books-validation
v3-small-dolma_common_crawl-validation
v3-small-dolma_pes2o-validation
v3-small-dolma_reddit-validation
v3-small-dolma_stack-validation
v3-small-dolma_wiki-validation
v3-small-ice-validation
v3-small-m2d2_s2orc-validation
v3-small-pile-validation
v3-small-wikitext_103-validation
Downstream Benchmarks
piqa (Bisk et al., 2020)
hellaswag (Zellers et al., 2019)
winogrande (Sakaguchi et al., 2020)
openbook_qa (Mihaylov et al., 2018)
sciq (Welbl et al., 2017)
arc_easy (Clark et al., 2018)
copa (Roemmele et al., 2011)
commitment_bank (De Marneffe et al., 2019)
mrpc (Dolan et al., 2005)
rte (Dagan et al., 2005)
sst2 (Socher et al., 2013)
G.4 Detailed Setup of Case-based and Rule-based Reasoning.
Following the work (Hu et al., 2024), we focus on binary operations that take two numbers, $a$ and $b$, as inputs. Denoting $c$ as the target label, the constructed datasets are in the form of $\mathcal{D}=\{((a_{i},b_{i}),c_{i})\}$ for two mathematical tasks: modular addition and linear regression. The two tasks are defined as follows:
  • Modular addition. The input to the model is “$a+b=$”, and the output is $c$, where $c=(a+b)\mod P$. The values of $a$ and $b$ range from 0 to 112. The constant $P$ is $113$ here.
  • Linear regression. This task involves the model learning a linear regression function. The input is given by “$(a,b)=$”, and the output is $c$, where $c=m\cdot a+n\cdot b+p$. The values of $a$ and $b$ range from 0 to 99. The constants are set as $m=1$, $n=2$, and $p=3$.
Leave-Square-Out
The work (Hu et al., 2024) employs the Leave-Square-Out method to evaluate the generalization ability of the Transformer. In this approach, a square test set is created to isolate the test samples from the training samples. For instance, consider the center of the square at $(a_{k},b_{k})$ with a side length of $l_{k}$. The square test set is defined as $\mathcal{T}_{k}=\{((a_{i},b_{i}),c_{i})\mid a_{k}-\frac{l_{k}}{2}\leq a_{i}\leq a_{k}+\frac{l_{k}}{2},b_{k}-\frac{l_{k}}{2}\leq b_{i}\leq b_{k}+\frac{l_{k}}{2}\}$, and all remaining samples from the training set. This division creates a "hole" in the center of the training set, which is more challenging for the model compared to a random split. Since there are no similar cases in the training set to aid the model in solving the problem, this method tests whether the model has truly learned the underlying rules. In the experiments of the work  (Hu et al., 2024), they found that Transformer-based models fail to generate correct answers for the test set in the "hole". Therefore, we use this method to assess the generalization ability of FANformer.
Settings
We finetune both the Transformer and FANformer models on each dataset for 500 epochs. The batch size is set to 336, and the learning rate is initialized at $10^{-4}$. A warm-up ratio of 0.01 is used, and we apply cosine decay to adjust the learning rate throughout the training process.
During generation, we set the model temperature to 0.5 and sample 10 generations to evaluate the accuracy on each test point. The square center $(a_{k},b_{k})$ is (50, 50) for linear regression and (56, 56) for modular addition.
Following the work (Hu et al., 2024), we apply the Leave-Square-Out method to each dataset. Specifically, we extract a square comprising 441 samples (from a total of approximately 10,000 samples) with a side length of 20 to form our test set, leaving the remainder as the training set. It is important to note that, despite removing a small portion of training samples, we ensure that all tokens present in the dataset appear in the training set. This precaution is to prevent the models from failing simply due to encountering unseen tokens. We then proceed to finetune Transformer and FANformer models using this specific training-test split for each dataset.